# SEVEN: Pruning Transformer Model by Reserving Sentinels

### v1

@@ II RELATED WORK
Pruning problems are primarily addressed using two major categories of methods [21]: unstructured pruning [22][23][24][11] and structured pruning [25][26][27][28]. Unstructured pruning involves scoring individual weights to remove single parameters from the entire model, while structured pruning involves removing complete modules from the model architecture, such as entire attention heads and layers. Additionally, some mask learning methods directly learn pruning masks during training and use them to prune weights [29]. Theoretical support for model pruning before training is provided by [30][8]. Building on this foundation, SNIP [9] and GraSP [10] explore the potential of single-shot pruning, i.e., pruning only once. Other works [31][12] achieve high-performance lottery tickets by performing multiple pruning of the model before training. In TM, pre-pruning is primarily based on weight magnitudes [8][12]. This scoring method originates from the inherent properties of pre-trained models and does not carry any fine-tuning-related information for the current task, resulting in suboptimal performance. In TM, dynamic pruning, carried out concurrently with training, has become mainstream. Movement [13], using first-order derivatives as scoring criteria, achieves good subnetworks at high pruning rates. It proposes a soft version, Soft-Movement, by suppressing the growth of the scoring function. PLATON [11] integrates the uncertainty of the scoring criteria to consider redundant weights. However, retaining weights with uncertainty affects the performance of subsequent fine-tuning. In contrast, the aim is to preserve weights that exhibit more stable performance. Before our work, Movement and PLATON methods provided the best performance for unstructured dynamic pruning.
@@ IV SEVEN
Where $\alpha_{1}$,$\alpha_{2}$ are hyperparameters in the range $\left(0,1\right)$, and $\overline{g_{0}}=\overline{g_{0}^{2}}=0$. A small subset of gradients from the original data is selected, denoted as $S=\{g_{1},g_{2},\ldots,g_{n}\}$, where $n<N$. Additionally, unbiased correction is applied to $\bar{g_{i}}$ and $\bar{g_{i}^{2}}$:
@@ VI Conclusion
MNLI RTE QNLI MRPC QQP SST-2 CoLA STS-B WNLI
$K$ 150 77 100 100 150 100 100 100 20
$\alpha_{1}$ 0.1 0.8 0.1 0.8 0.1 0.5 0.8 0.8 0.8
$\alpha_{2}$ 0.2 0.9 0.2 0.9 0.1 0.7 0.9 0.9 0.9
@@
MNLI RTE QNLI MRPC QQP SST-2 CoLA STS-B
$t_{i}$ 5400 200 2000 300 5400 1000 500 500
$K$ 16600 1000 10000 600 16600 4000 1000 2000
$\alpha_{1}$ 0.7 0.8 0.8 0.7 0.7 0.8 0.8 0.8
$\alpha_{2}$ 0.9 0.8 0.9 0.9 0.9 0.8 0.8 0.8

### v2

@@ II RELATED WORK
Pruning problems are primarily addressed using two major categories of methods [21]: unstructured pruning [22][23][24][11] and structured pruning [25][26][27][28]. Unstructured pruning involves scoring individual weights to remove single parameters from the entire model, while structured pruning involves removing complete modules from the model architecture, such as entire attention heads and layers. Additionally, some mask learning methods directly learn pruning masks during training and use them to prune weights [29]. Theoretical support for model pruning before training is provided by [30][8]. Building on this foundation, SNIP [9] and GraSP [10] explore the potential of single-shot pruning, i.e., pruning only once. Other works [31][12][32] achieve high-performance lottery tickets by performing multiple pruning of the model before training. In TM, pre-pruning is primarily based on weight magnitudes [8][12]. This scoring method originates from the inherent properties of pre-trained models and does not carry any fine-tuning-related information for the current task, resulting in suboptimal performance. In TM, dynamic pruning, carried out concurrently with training, has become mainstream. Movement [13], using first-order derivatives as scoring criteria, achieves good subnetworks at high pruning rates. It proposes a soft version, Soft-Movement, by suppressing the growth of the scoring function. PLATON [11] integrates the uncertainty of the scoring criteria to consider redundant weights. However, retaining weights with uncertainty affects the performance of subsequent fine-tuning. In contrast, the aim is to preserve weights that exhibit more stable performance. Before our work, Movement and PLATON methods provided the best performance for unstructured dynamic pruning.
@@ IV SEVEN
Where $\alpha_{1}$$\alpha_{2}$ are hyperparameters in the range $\left(0,1\right)$, and $\overline{g_{0}}=\overline{g_{0}^{2}}=0$. A small subset of gradients from the original data is selected, denoted as $S=\{g_{1},g_{2},\ldots,g_{n}\}$, where $n<N$. Additionally, unbiased correction is applied to $\bar{g_{i}}$ and $\bar{g_{i}^{2}}$:
@@ VI Conclusion
MNLI RTE QNLI MRPC QQP
$K$ 150 77 100 100 150
$\alpha_{1}$ 0.1 0.8 0.1 0.8 0.1
$\alpha_{2}$ 0.2 0.9 0.2 0.9 0.1
SST-2 CoLA STS-B WNLI
$K$ 100 100 100 20
$\alpha_{1}$ 0.5 0.8 0.8 0.8
$\alpha_{2}$ 0.7 0.9 0.9 0.9
@@
MNLI RTE QNLI MRPC
$t_{i}$ 5400 200 2000 300
$K$ 16600 1000 10000 600
$\alpha_{1}$ 0.7 0.8 0.8 0.7
$\alpha_{2}$ 0.9 0.8 0.9 0.9
QQP SST-2 CoLA STS-B
$t_{i}$ 5400 1000 500 500
$K$ 16600 4000 1000 2000
$\alpha_{1}$ 0.7 0.8 0.8 0.8
$\alpha_{2}$ 0.9 0.8 0.8 0.8